Grouper Company changed depreciation methods in 2017 from double-declining-balance to straight-line. Depreciation prior to 2017 under double-declining-balance was $97,100, whereas straight-line depreciation prior to 2017 would have been $49,600. Grouper’s depreciable assets had a cost of $254,300 with a $42,000 salvage value, and an 8-year remaining useful life at the beginning of 2017.

Prepare the 2017 journal entry related to Grouper’s depreciable assets (Equipment). (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Account Titles and Explanation

Debit

Credit
Answer #1
Account titles and explanation Debit Credit
Depreciation $ 26,538
To Equipment $26,538

Calculation of Depreciation :

Depreciation = Original cost - Slavage Value / Life of the asset = $ 2,54,300 - $ 42,000/8

=$ 26,538
Answer #2

254,300 - 97,100 = 157,200 - 42000 = 115,200

115,200/8 = 14,400


Debit: Depreciation 14,400

Credit: Accumulated Depreciation - Equipment 14,400